OpenStudy (anonymous):

find the 12th term of each arithmetic sequence a4= 88 & a8=29

OpenStudy (anonymous):

a12 = 29 + (88-29)

OpenStudy (anonymous):

whaat?

OpenStudy (anonymous):

@ankit042

OpenStudy (ankit042):

a(12) = a(0) + (12-1)d where you have to first find value of a(0) and d

OpenStudy (ankit042):

use a(4)= a(0) +(4-1)d and a(8) = a(0) +(8-1)d

OpenStudy (anonymous):

... you just lost me ;0 i dont understand this at all...

OpenStudy (reemii):

from a4 to a5, a certain quantity X is added to a4. from a5 to a6, a certain quantity X is added to a5. etc. so that a4 + X + X + X + X = a8 to compute a12, you must add X, X, X, and X. (4X) It turns out that if you look at a8-a4, it is already equal to 4X, so you don't have to know X personally, just 4X.

OpenStudy (reemii):

(* to compute a12, add 4X to a8)

OpenStudy (reemii):

what is 12? a4 = 88 a8 = 29. Therefore, 4X = 29 - 88 = -59, and a12 = a8 + 4X = 29 + (-59).

OpenStudy (reemii):

ok. ( arithmetic sequence: from one number to another one, the SAME quantity is added each time. that's what is noted 'X' in my answers, and 'd' in @ankit042 's one. To compute a12, we look at the difference between a8 and a4 so get an idea of what that quantity 'X' is. re-read the explanations. (btw the ans. is -30)

OpenStudy (anonymous):

Thanks :/

OpenStudy (reemii):

(look at this simpler example.. a4 = 5, a5 = 6. find a6. you see that from a4 to a5, the number 1 is added. this is the 'X' (or 'd'). a6 is simply a5 + 1. your question is similar, but with other numbers.)

OpenStudy (anonymous):

ohhh i think i get it nooow. do you think you could help with another?

OpenStudy (reemii):

ok type your question

OpenStudy (ankit042):

a new question you mean :)

OpenStudy (anonymous):

a20= 1.4 & a30 = 7.2

OpenStudy (reemii):

what do you have to compute? which a... ?

OpenStudy (anonymous):

the 12th term , saame as before.

OpenStudy (reemii):

ok. the first thing is to find the 'X' between two consecutive numbers. Between a20=... and a30=... , the difference is not 'X' but 10'X' because a30 is 10 steps further in the sequence. ok?

OpenStudy (anonymous):

ohkay (:

OpenStudy (reemii):

so what is X ? do you have to add or remove a few X's from a20 to find a12 ?

OpenStudy (anonymous):

add ?

OpenStudy (reemii):



OpenStudy (anonymous):

x from a 20 to find

OpenStudy (anonymous):

oh yeaah , ohkay ...

OpenStudy (reemii):

find X, and then you find a12 = a20 - 8X = 1.4 - 8 X = ...

OpenStudy (reemii):

I explained above how to find X. ——— Between a20=... and a30=... , the difference is not 'X' but 10'X' because a30 is 10 steps further in the sequence. ——— a30 - a20 = 10X

OpenStudy (anonymous):

-8.6 ?

OpenStudy (reemii):

a30 = 7.2, a20 = 1.4 a30-a20 = 7.2 - 1.4 = is not -8.6

OpenStudy (anonymous):

did u find it out

OpenStudy (anonymous):

ohkay nevermind i dont get it , i give up ;0 im tired of math.

OpenStudy (reemii):

call it a break :-|

OpenStudy (anonymous):

alright..